Find the WACC, assume that companys tax rate is 40% Common stock: 15 million shares outstanding, selling for 5$ per share; thebeta is 1.05 Preffered Stock:5 million snares cutstanding, sallingfor $4.5 per share, pays $0.9 annually per share. Debt 1 million 8% quarters coupon bonds outstanding, $100 face value, 15 years to maturity, selling at Market: 6.5% market return and 4.5% risk-freerate.
Answer #1
Cost of common equity per CAPM = risk free rate+beta*(market return-risk free rate) =
= 4.5%+1.05*(6.5%-4.5%) = 6.60%
Cost of preferred stock = Dividend/Price = 0.9/4.5 = 20.00%
Cost of debt = 8%*(1-40%) = 4.80%
CALCULATION OF WACC:
Component Market Value           ($ million) Weight Component Cost WACC
Common equity [15*5] 75.00 37.97% 6.60% 2.51%
Preferred stock [5*4.5] 22.50 11.39% 20.00% 2.28%
Debt [1*100] 100.00 50.63% 4.80% 2.43%
197.50 7.22%
WACC = 7.22%
Note:
*The cost of common stock seems to be too low compared to the cost of preferred stock.
*The cost of debt is taken as 8% (before tax). Stated as 8% quarters is not clear.